OpenStudy (anonymous):

which of the following ratios are equivalent 5/20? Choose all that apply. A. 1/4 B. 1/5 C. 20/100 D. 25/100

OpenStudy (solomonzelman):

try to divide top and bottom by 5. 5/20= (5/5) / (20/5) = ?

OpenStudy (anonymous):

(5/5=1) / (20/5=4)

OpenStudy (anonymous):

4/1=4

OpenStudy (solomonzelman):

No,you did the 1st step correctly, but not the second you have, 5/20= (5/5) / (20/5) = 1 /4 and that's not equal to 4, rather - it is which of the choices? See?

OpenStudy (anonymous):

Yes, thx

OpenStudy (solomonzelman):

OK, YW!